How would you solve for the slope?
(-9,6) and (-4,5)

Answer by TP(29)

slope(gradient)of straight line between two points on the line is vertical distance between the two points divided by the horizontal distance between the two points.
So if the two points on the line have coordinates of (X,Y) and (x,y) then
the slope=(Y-y)/(X-x).
In your question the two points on the line are (-9,6) and (-4,5) and so here
X=-9,x=-4 and Y=6,y=5.
Replacing the letters X,x,Y and y by their number values the formula becomes
the slope=(6-5)/(-9--4).
Now the top part=1 and the bottom part =-9--4 becomes -9+4 (because --=+).
The value of -9+4 is -5 (by using the number line,start at -9 and move 4 spaces to the right and you end up at -5).
So the slope is 1/-5=-0.2ANS